In an isosceles triangle, the length of one of the equal sides is three time the length of the third side. The perimeter is 21 meters. Find the length of each side.
Answer by checkley77(12844)

3s+3s+s=21
7s=21
s=3 For the third short side.
3*3=9 for the equal sides.
Proof:
9+9+3=21
21=21